OpenStudy (anonymous):

please ~ A line has an x-intercept of (2,0) and a y-intercept of (0, -4) Determine the equation of the line in Standard Form.

OpenStudy (mathmate):

In symmetric form it is x/2 + y/(-4)=1 Multiply to get rid of fractions, 2x-y=4 Standard form y=2x-4 General form 2x-y-4=0

OpenStudy (tkhunny):

Also called Intercept Form! I'm delighted. I see it so rarely. \(\dfrac{x}{2}+\dfrac{y}{-4} = 1\)